# ADS-Edit: A Multimodal Knowledge Editing Dataset for Autonomous Driving Systems

### v1

@@
Abstract
Recent advancements in Large Multimodal Models (LMMs) have shown promise in Autonomous Driving Systems (ADS). However, their direct application to ADS is hindered by challenges such as misunderstanding of traffic knowledge, complex road conditions, and diverse states of vehicle. To address these challenges, we propose the use of Knowledge Editing, which enables targeted modifications to a model’s behavior without the need for full retraining. Meanwhile, we introduce ADS-Edit, a multimodal knowledge editing dataset specifically designed for ADS, which includes various real-world scenarios, multiple data types, and comprehensive evaluation metrics. We conduct comprehensive experiments and derive several interesting conclusions. We hope that our work will contribute to the further advancement of knowledge editing applications in the field of autonomous driving111Code and data are available in https://github.com/zjunlp/EasyEdit..
ADS-Edit: A Multimodal Knowledge Editing Dataset for
Autonomous Driving Systems
Chenxi Wang♡11footnotemark: 1, Jizhan Fang♡11footnotemark: 1, Xiang Chen♠††thanks: Equal Contribution., Bozhong Tian♡, Ziwen Xu♡, Huajun Chen♡, Ningyu Zhang♡††thanks: Corresponding author. ♡Zhejiang University ♠Nanjing University of Aeronautics and Astronautics {sunnywcx,zhangningyu}@zju.edu.cn
1 Introduction
The recent Large Multimodal Models (LMMs) (Wang et al., 2024a; Li et al., 2024a; Wu et al., 2024b; Liu et al., 2024c; Team, 2024, 2025) have significantly enhanced capabilities in video understanding and multimodal reasoning. As a key foundation, LMMs have also found initial applications in Autonomous Driving Systems (ADS) (Xing et al., 2024; Sima et al., 2024; Huang et al., 2024b; Yao et al., 2024). However, the direct application of LMMs in ADS yields suboptimal results.
@@ 1 Introduction
To address these challenges, we propose the use of Knowledge Editing, which enables targeted modifications to the model’s behavior without the computational burden of full retraining. Unlike traditional fine-tuning, which risks catastrophic forgetting and demands extensive resources, Knowledge Editing facilitate rapid knowledge updates by selectively altering parameters associated with specific factual or contextual knowledge (Yao et al., 2023; Gupta et al., 2024; Zhang et al., 2024d; Fang et al., 2024a; Youssef et al., 2025; Jiang et al., 2025; Liu et al., 2024a; Yao et al., 2025). Although Knowledge Editing has been preliminarily explored in the area of LMMs, existing efforts have primarily focused on editing multimodal common knowledge (Cheng et al., 2023; Ma et al., 2024) and multimodal factual knowledge (Zhang et al., 2024a; Huang et al., 2024a). To address this limitation, we pioneer the editing of multimodal domain-specific knowledge in the field of autonomous driving.
Following the principle of leveraging knowledge editing to address the challenges of autonomous driving, we design the ADS-Edit benchmark. This benchmark encompasses three real-world scenarios: perception, understanding, and decision making. Additionally, it incorporates three types of data: video, multi-view images, and single image. Furthermore, we establish comprehensive evaluation metrics for knowledge editing in autonomous driving scenarios.
We evaluate four commonly used Knowledge Editing baselines under both single editing and lifelong editing scenarios, such as Prompt, AdaLora (Zhang et al., 2023), GRACE (Hartvigsen et al., 2023), and WISE (Wang et al., 2024b). Thirdly, through our analysis, we obtained a series of interesting findings, including the universality of knowledge editing methods in updating knowledge across various scenarios, their ability to balance editing effectiveness and processing speed, particularly in the context of video data, and the remaining limitations in locality. In general, we conclude our contributions as:
@@ 2 Background
We propose leveraging knowledge editing techniques to address challenges in autonomous driving scenarios, defined as follows: Let be a LMM and be an edited model. Given the user’s inputs (Includes text and autonomous driving’s multimodal inputs , such as image or videos) and the editing target , the edited model is expected to modify LMM’s outputs within the editing scope to match the editing target, while preserving the original model’s output for the inputs outside the editing scope. The specific definition is as follows:
where and means the in-scope of the editing inputs.
3 Bechmark Construction
3.1 Desgin Principle
@@ 3.2 Data Collection
3.2.1 Reliability Data Construction
We select three well-known datasets of the autonomous driving system: LingoQA (Marcu et al., 2024), DriveLM dataset (Sima et al., 2024), and CODA-LM dataset (Li et al., 2024d) as raw data. Their autonomous driving scenarios consist of video (5 frames), multi-view images (6 views), and single-image, respectively.
Data Preprocess.
Notably, the answers in the raw data consist of more tokens (e.g., detailed driving suggestions), distinguishing them from unimodal knowledge editing tasks, such as those represented by triple-based, which have concise token answers. This introduces the following challenges: 1) Suboptimal Editing Performance. Redundant answers pose difficulties for the implementation of certain editing methods, particularly causal tracing-based approaches such as ROME (Meng et al., 2022). 2) Difficult Evaluation. Evaluating the accuracy of long text sequences often requires external models to score semantic consistency, which not only suffers from limited accuracy but also incurs high computational costs. Consequently, simplifying the answers in the raw data becomes a necessary step, as shown in Figure 3. We prompt Deepseek-v3 (et al, 2024) to condense the original answers (The prompt template is shown in Appendix C). Meanwhile, we retain the original answers as reserved information for potential use in future research.
Reliability Data Construction.
We observe that the data from DriveLM include predictions of vehicle trajectories, which are not feasible for streamlined and generalized data collection. Consequently, we decide to exclude it. For CODA-LM, which contains a significant amount of autonomous driving suggestions and image descriptions, we opt to use Deepseek-v3 to self-generate QA pairs (The prompt template is shown in Appendix C). Subsequently, we shuffle all the data and select a subset to serve as reliability data.
3.2.2 Generality Data Construction
@@
Given the same driving scenario, the text query is modified to test whether the model truly understands the underlying context. Following prior works, we prompt Deepseek-v3 to rewrite the original query into a semantically similar query while keeping the answer unchanged (See the prompt in Appendix C).
@@
3.2.3 Locality Data Construction
@@
To evaluate the capability of editing methods in preserving text locality, we select the Natural Questions dataset (Kwiatkowski et al., 2019), which contains unimodal factual and commonsense knowledge, and randomly sample a subset of it to serve as Textual Locality data. For assessing the preservation of multimodal general capabilities, we select a portion of the VQAv2 dataset (Antol et al., 2015), which includes general visual question answers, to construct Multimodal Locality data.
| Method | Reliability | T-Generality | M-Generality | T-Locality | M-Locality | |
| LLaVA-Onevision | ||||||
| Prompt | 94.25 | 90.18 | 95.04 | 84.86 | 68.13 | |
| AdaLora | 78.01 | 72.76 | 75.84 | 85.51 | 81.12 | |
| GRACE | 100.00 | 28.91 | 28.16 | 100.00 | 100.00 | |
| WISE | 99.10 | 86.97 | 95.78 | 94.18 | 99.98 | |
| Qwen2-VL | ||||||
| Prompt | 90.57 | 84.98 | 90.48 | 89.61 | 72.44 | |
| AdaLora | 79.89 | 75.68 | 78.76 | 82.27 | 69.37 | |
| GRACE | 100.00 | 27.01 | 29.93 | 100.00 | 100.00 | |
| WISE | 94.18 | 85.20 | 91.99 | 94.23 | 99.85 | |
3.3 Quality Control
@@
As outlined in the aforementioned process and illustrated in Figure 3, the entire pipeline is fully automated. However, ensuring data quality through manual verification remains a critical step. To this end, we conduct manual verification and calibration for all processes involving AI models, including answer simplification, QA pairs generation from CODA-LM, textual generality data generation, and scenario categorization. Furthermore, for multimodal generality data, we manually sample 20% of the data to assess similarity and filter out those instances with no resemblance. The data quality control tasks are evenly distributed among three annotators with graduate-level education, who independently performed the calibration. In cases of uncertainty, the annotators first independently judge on whether modifications are necessary, and then resolve any disagreements through a majority vote. The inter-annotator consistency, measured by Fleiss’s Kappa (), demonstrates substantial agreement ( = 0.912), based on a randomly selected sample of 200 annotated data points, indicating a high level of consistency within the range of 0.80 ≤ ≤ 1.00.
4 Evaluation
4.1 Experimental settings
@@
We conduct experiments on the most advanced LMMs to date, which are LLaVA-OneVision Li et al. (2024a) and Qwen2-VL Wang et al. (2024a). Both models utilize Qwen2-7B as their LLM component. Meanwhile, four classical knowledge editing baselines, including Prompt, AdaLora, GRACE, and WISE are employed to update the knowledge of LMMs. Further details can be found in the Appendix B. To comprehensively evaluate the performance of various baselines in ADS-Edit, we test in Single Editing and Lifelong Editing scenarios. Single editing involves updating and evaluating the model immediately after receiving each individual driving data instance. In contrast, lifelong editing represents a more realistic driving scenario, where multiple driving data instances are collected during the vehicle’s moving, requiring continuous integration of this knowledge into the model followed by evaluation. We reasonably assume that the visual processing component (e.g., Vision Transformer) of LMMs provides reliable visual information. Therefore, we focus only on editing the LLM component to modify its understanding of visual input.
4.2 Main Results
4.2.1 Single Editing Results
Memory-based editing methods, such as GRACE and WISE, have proven highly effective in modulating the behavior of LMMs for autonomous driving task prediction. Notably, GRACE achieves a 100% modification rate in both LLaVA-Onevision and Qwen2-VL. Although the ground truth answer is directly provided in the input text, the reliability of the Prompt remains suboptimal. We hypothesize that this is due to the model being affected by interference when processing multimodal driving inputs, which differs from the unimodal scenarios. Under a limited number of training epochs, AdaLora struggles to achieve satisfactory reliability through model parameter updates. Evaluating the generalization of editing methods to other autonomous driving scenarios and queries is essential. However, GRACE exhibits the worst prediction performance on Generality, with accuracy dropping below 30% in both LMMs, compared to its Reliability. This is attributed to the difficulty of GRACE’s codebook in capturing the representational differences of long-sequence multimodal inputs. The remaining three baselines show similar Generality and Reliability results. Among them, WISE and Prompt achieve approximately 90% on Generality, demonstrating their stable generalization capabilities. AdaLora still struggle to generalize the other driving scenarios. Furthermore, M-Generality of the baselines exceeds T-Generality, suggesting that LMMs tend to focus more on text tokens while making less use of visual tokens. This observation may inspire further research on enhancing the efficient utilization of redundant visual representations in LMMs. While preserving the model’s local capabilities, GRACE demonstrates a significant advantage in Locality. WISE, which employs a dual-memory mechanism, also maintains good Locality, achieving nearly 100%. In contrast, AdaLora and Prompt somewhat disrupt the original model’s Locality. Additionally, we observe that both AdaLora and Prompt exhibit lower M-Locality than T-Locality, which is due to the changes in the LMM’s behavior caused by the multimodal inputs in autonomous driving scenarios.
4.2.2 Lifelong Editing Results
4.3 Analysis
@@
Although both models use Qwen2 as Large Language Model, LLaVA-Onevision achieve better results across most metrics in terms of reliability and generality. Specifically, on Prompt and WISE methods, Qwen2-VL demonstrates stronger performance in retaining the original predicted answers. Qwen2-VL slightly outperforms LLaVA-OneVision in terms of reliability and generality by using AdaLora. However, locality is significantly compromised compared to LLaVA-OneVision.
@@
5 Related Work
5.1 Knowledge Editing
@@
Recent advances in knowledge editing have emerged as a pivotal research direction in knowledge updating (Chen, 2024; Liu et al., 2024b; Wang et al., 2024c; Wu et al., 2024a; Xu et al., 2024; Li and Chu, 2024; Nie et al., 2024; Wei et al., 2025; Zhang et al., 2025; Ali et al., 2025; Hwang et al., 2025; Zhao et al., 2025; Markowitz et al., 2025; Yang et al., 2025; Dong et al., 2025; Pan et al., 2025; Wang et al., 2025; Liu et al., ; Zhang et al., 2024b; Xu et al., 2025; Bi et al., 2025; He et al., 2025; Li et al., 2025; Shen and Huang, 2025). Following Yao et al. (2023), we systematize current methodologies into two principal paradigms: 1) Parameter Modification Approaches: These methods directly alter the model’s internal representations through targeted interventions (Meng et al., 2022, 2023; Fang et al., 2024b; Cai and Cao, 2024; Mitchell et al., 2022; Zhang et al., 2024c; Tan et al., 2024; Hu et al., 2022; Zhang et al., 2023; Dettmers et al., 2023; Feng et al., 2025). 2) Parameter Preservation Approaches: These strategies maintain original parameters while updating new knowledge (Zheng et al., 2023; Jiang et al., 2024; Wang et al., 2024b; Hartvigsen et al., 2023; Li et al., 2025). The emergence of Large Multimodal Models (LMMs) has also advanced the field of multimodal knowledge editing. These works mainly focus on the editing of multimodal common knowledge (Cheng et al., 2023; Ma et al., 2024; Pan et al., 2024; Wang et al., 2024d) or multimodal factual knowledge (Li et al., 2024b; Zhang et al., 2024a; Huang et al., 2024a). However, the application of multimodal knowledge editing in the context of domain-specific knowledge, like autonomous driving remains underexplored.
5.2 Large Multi-modal Models for Autonomous Driving

### v2

@@
Abstract.
Recent advancements in Large Multimodal Models (LMMs) have shown promise in Autonomous Driving Systems (ADS). However, their direct application to ADS is hindered by challenges such as misunderstanding of traffic knowledge, complex road conditions, and diverse states of vehicle. To address these challenges, we propose the use of Knowledge Editing, which enables targeted modifications to a model’s behavior without the need for full retraining. Meanwhile, we introduce ADS-Edit, a multimodal knowledge editing dataset specifically designed for ADS, which includes various real-world scenarios, multiple data types, and comprehensive evaluation metrics. We conduct comprehensive experiments and derive several interesting conclusions. We hope that our work will contribute to the further advancement of knowledge editing applications in the field of autonomous driving111Code and data are available in https://github.com/zjunlp/EasyEdit/blob/main/examples/ADSEdit.md..
1. Introduction
The recent Large Multimodal Models (LMMs) (Wang et al., 2024a; Li et al., 2024d; Wu et al., 2024a; Liu et al., 2024c; Team, 2024, 2025) have significantly enhanced capabilities in video understanding and multimodal reasoning. As a key foundation, LMMs have also found initial applications in Autonomous Driving Systems (ADS) (Xing et al., 2024; Sima et al., 2024; Huang et al., 2024a; Yao et al., 2024). However, the direct application of LMMs in ADS yields suboptimal results.
@@ 1. Introduction
To address these challenges, we propose the use of Knowledge Editing, which enables targeted modifications to the model’s behavior, eschewing the substantial computational demands associated with full retraining. Unlike traditional fine-tuning, which risks catastrophic forgetting and demands extensive resources, Knowledge Editing facilitate rapid knowledge updates by selectively altering parameters associated with specific factual or contextual knowledge (Yao et al., 2023; Gupta et al., 2024; Zhang et al., 2024b; Fang et al., 2024a; Youssef et al., 2025; Jiang et al., 2025; Liu et al., 2024a; Yao et al., 2025). While preliminary explorations of Knowledge Editing in the area of LMMs have predominantly concentrated on the manipulation of multimodal common knowledge (Cheng et al., 2023; Ma et al., 2024) and multimodal factual knowledge (Zhang et al., 2024d; Huang et al., 2024b), we pioneer the editing of multimodal domain-specific knowledge in the field of autonomous driving.
Following the principle of leveraging knowledge editing to address the challenges of autonomous driving system, we propose the ADS-Edit benchmark. This benchmark encompasses three real-world scenarios: perception, understanding, and decision making. Additionally, it incorporates three types of data: video, multi-view images, and single image. Furthermore, we establish comprehensive evaluation metrics for knowledge editing in autonomous driving scenarios.
We evaluate four commonly used Knowledge Editing baselines under both single editing and lifelong editing scenarios, such as Prompt, AdaLora (Zhang et al., 2023), GRACE (Hartvigsen et al., 2023), and WISE (Wang et al., 2024d). Through our analysis, we obtained a series of interesting findings, including the universality of knowledge editing methods in updating knowledge across various scenarios, their ability to balance editing effectiveness and processing speed, particularly in the context of video data, and the remaining limitations in locality. Moreover, in real-world evaluation scenarios, it is evident that current editing techniques require further refinement and enhancement to meet the demands of autonomous driving systems.
In general, we conclude our contributions as:
@@ 2. Background
We propose leveraging knowledge editing techniques to address the challenges encountered in autonomous driving scenarios. Formally, let denote a multimodal large model (LMM), and let represent the edited model. Given a user input , consisting of textual input and multimodal signals from autonomous driving (e.g., images or videos), along with a desired editing target , the goal is to update the model such that . The edited model should modify the output of the original LMM within the intended editing scope while preserving its original behavior on inputs outside that scope. The formal definition is as follows:
where and means in-scope of editing inputs.
3. Bechmark Construction
3.1. Desgin Principle
@@ 3.2. Data Collection
3.2.1. Data Preprocess.
We select three well-known datasets of the autonomous driving system: LingoQA (Marcu et al., 2024), DriveLM dataset (Sima et al., 2024), and CODA-LM dataset (Li et al., 2024b) as raw data. These datasets encompass a variety of autonomous driving scenarios, including video (5 frames), multi-view images (6 perspectives), and single images, respectively. Notably, the answers in the raw data consist of more tokens (e.g., detailed driving suggestions), distinguishing them from unimodal knowledge editing tasks, such as those represented by triple-based formats, which typically feature concise token responses. This introduces the following challenges: 1) Suboptimal Editing Performance. Redundant answers pose difficulties for the implementation of certain editing methods, particularly causal tracing-based approaches such as ROME (Meng et al., 2022). 2) Difficult Evaluation. Evaluating the accuracy of long text sequences often requires external models to score semantic consistency, which not only suffers from limited accuracy but also incurs high computational costs. Consequently, simplifying the answers in the raw data becomes a necessary step, as shown in Figure 3. We prompt Deepseek-v3 (et al, 2024) to condense the original answers.Meanwhile, we retain the original answers as meta information for potential use in future research.
3.2.2. Reliability Data Construction
We observe that the data from DriveLM include predictions of vehicle trajectories, which are not feasible for streamlined and generalized data collection. Consequently, we decide to exclude it. For CODA-LM, which contains a significant amount of autonomous driving suggestions and image descriptions, we opt to use Deepseek-v3 to self-generate QA pairs.We then shuffle all the data and select a subset to serve as reliability data. Furthermore, to present the metadata from multiple perspectives, we categorize the selected data by scenario types, comprising three main categories and eleven subcategories. Specifically, we prompt DeepSeek-v3 to perform classification, which is then verified through rigorous manual inspection, as detailed in Section 3.3.
Method | Reliability | T-Generality | M-Generality | T-Locality | M-Locality | ||||||
Edit. | Real. | Edit. | Real. | Edit. | Real. | Edit. | Real. | Edit. | Real. | ||
LLaVA-Onevision | |||||||||||
Prompt | 94.25 | 97.68 | 90.18 | 94.71 | 95.04 | 97.21 | 84.47 | 11.87 | 80.35 | 79.13 | |
AdaLora | 78.01 | 56.68 | 72.76 | 54.36 | 75.84 | 53.34 | 85.51 | 40.54 | 81.12 | 69.67 | |
GRACE | 100.00 | 89.52 | 28.91 | 45.83 | 28.16 | 45.45 | 100.00 | 100.00 | 100.00 | 100.00 | |
WISE | 99.10 | 75.97 | 86.97 | 70.87 | 95.78 | 73.56 | 94.18 | 56.96 | 99.98 | 82.84 | |
Qwen2-VL | |||||||||||
Prompt | 90.57 | 97.77 | 84.98 | 94.53 | 90.48 | 97.03 | 89.61 | 61.04 | 72.44 | 66.14 | |
AdaLora | 79.89 | 65.03 | 75.68 | 61.69 | 78.76 | 63.64 | 82.27 | 41.37 | 69.37 | 57.33 | |
GRACE | 100.00 | 51.67 | 27.01 | 47.96 | 29.93 | 52.41 | 100.00 | 100.00 | 100.00 | 100.00 | |
WISE | 94.18 | 62.98 | 85.20 | 57.14 | 91.99 | 62.15 | 94.23 | 70.13 | 99.85 | 71.33 |
3.2.3. Generality Data Construction
@@
Given the same driving scenario, the text query is modified to test whether the model truly understands the underlying context. Following prior works, we prompt Deepseek-v3 to rewrite the original query into a semantically similar query while keeping the answer unchanged.
@@
3.2.4. Locality Data Construction
@@
To evaluate the ability of editing methods to preserve textual locality, we utilize the Natural Questions dataset (Kwiatkowski et al., 2019), which encompasses unimodal factual and commonsense knowledge. A random subset of this dataset is sampled to constitute the Textual Locality data. For assessing the preservation of general multimodal capabilities, we select a portion of the VQAv2 dataset (Antol et al., 2015), comprising general visual question-answer pairs, to form the Multimodal Locality data.
3.3. Quality Control
@@
As outlined in the aforementioned process and illustrated in Figure 3, the entire pipeline is fully automated. However, ensuring data quality through manual verification remains a critical step. To this end, we conduct manual verification and calibration for all processes involving AI models, including answer simplification, QA pairs generation from CODA-LM, textual generality data generation, and scenario categorization. Furthermore, for multimodal generality data, we manually sample 20% of the data to assess similarity and filter out those instances with no resemblance. To ensure the consistency between the rationales in the generality data answers and the reliability data, we filter out mismatched samples by referring to the original answers. The data quality control tasks are evenly distributed among three annotators with graduate-level education, who independently performed the calibration. In cases of uncertainty, the annotators first independently judge on whether modifications are necessary, and then resolve any disagreements through a majority vote. The inter-annotator consistency, measured by Fleiss’s Kappa (), demonstrates substantial agreement ( = 0.802), based on a randomly selected sample of 200 annotated data points, indicating a high level of consistency within the range of 0.80 ≤ ≤ 1.00.
4. Evaluation
4.1. Experimental settings
@@
We conduct experiments on the most advanced LMMs to date, which are LLaVA-OneVision (Li et al., 2024d) and Qwen2-VL (Wang et al., 2024a). Both models utilize Qwen2-7B as their LLM component. Meanwhile, four classical knowledge editing baselines, including Prompt, AdaLora, GRACE, and WISE are employed to update the knowledge of LMMs. Our analysis covers both the Single Editing and Lifelong Editing settings to thoroughly assess the performance of different baselines in ADS-Edit. Single editing involves updating and evaluating the model immediately after receiving each individual driving data instance. In contrast, lifelong editing represents a more realistic driving scenario, where multiple driving data instances are collected during the vehicle’s moving, requiring continuous integration of this knowledge into the model followed by evaluation. We reasonably assume that the visual processing component (e.g., Vision Transformer) of LMMs provides reliable visual information. Therefore, we focus only on editing the LLM component to modify its understanding of visual input. To comprehensively evaluate the effectiveness of editing methods on ADS, we adopt two evaluation metrics: (1) Editing evaluation (Edit), where tokens are generated in a teacher-forcing manner and compared at the token level (Meng et al., 2022). (2) Real-world evaluation (Real), where the edited model performs free-form generation, and a LLM is prompted to assess the semantic consistency between the generated outputs and the ground-truth answers (Yang et al., 2025a). Due to cost and resource constraints, we report the Real metric only for the main single-edit experiment, while the remaining experiments are evaluated using the Edit metric by default.
4.2. Main Results
We present the results for the single-edit scenario below; the results for lifelong editing can be found on Github222https://github.com/zjunlp/EasyEdit/blob/main/examples/ADSEdit.md.
Results of Editing evaluation.
Memory-based editing methods, such as GRACE and WISE, have proven highly effective in modulating the behavior of LMMs for autonomous driving task prediction. Notably, GRACE achieves a 100% modification rate in both models, yet performs poorly in terms of generality. WISE demonstrates a well-balanced performance across reliability, generality, and locality. Although the ground truth answer is directly provided in the input text, the reliability and generality of the Prompt remains suboptimal. Meanwhile, Prompt performs poorly on the locality, which can be attributed to its sensitivity to the ground-truth answers in the input. Under a limited resources, AdaLora struggles to achieve satisfactory reliability and generality through model parameter updates. Moreover, parameter modifications significantly impact Adalora’s performance on the locality metric.
Results of Real-world evaluation.
Although Prompt underperforms GRACE and WISE on conventional editing metrics, it achieves the best performance in terms of reliability and generalization under real-world evaluation settings. This is primarily because the model does not strictly replicate the ground-truth answers but instead generates outputs that are semantically aligned with them. In contrast, the other three editing methods exhibit a noticeable decline in performance when transitioning from editing metrics to real-world evaluation, indicating limitations in their reliability and generalization for ADS applications. In terms of locality metrics, GRACE and WISE demonstrate strong performance, whereas Prompt underperforms.
4.3. Analysis
@@
Although both models use Qwen2 as Large Language Model, LLaVA-Onevision achieve better results across most metrics in terms of reliability and generality. Specifically, on Prompt and WISE methods, Qwen2-VL demonstrates stronger performance in retaining the original predicted answers. Qwen2-VL slightly outperforms LLaVA-OneVision in terms of reliability and generality by using AdaLora. However, locality is significantly compromised compared to LLaVA-OneVision. Furthermore, in real-world evaluation scores, we observe that Qwen2-VL tends to retain its original outputs and exhibits resistance to editing.
@@
5. Related Work
5.1. Knowledge Editing
@@
Recent advances in knowledge editing have emerged as a pivotal research direction in knowledge updating (Chen, 2024; Liu et al., 2024b; Wang et al., 2024b; Wu et al., 2024b; Xu et al., 2024; Li and Chu, 2024; Nie et al., 2024; Wei et al., 2025; Zhang et al., 2025; Ali et al., 2025; Hwang et al., 2025; Zhao et al., 2025; Markowitz et al., 2025; Yang et al., 2025b; Dong et al., 2025; Pan et al., 2025; Wang et al., 2025; Liu et al., [n. d.]; Zhang et al., 2024c; Xu et al., 2025; Bi et al., 2025; He et al., 2025; Li et al., 2025; Shen and Huang, 2025; Yang et al., 2025a). Following Yao et al. (2023), we systematize current methodologies into two principal paradigms: (1) Parameter Modification Approaches: These methods directly alter the model’s internal representations through targeted interventions (Meng et al., 2022, 2023; Fang et al., 2024b; Cai and Cao, 2024; Mitchell et al., 2022; Zhang et al., 2024a; Tan et al., 2024; Hu et al., 2022; Zhang et al., 2023; Dettmers et al., 2023; Feng et al., 2025). (2) Parameter Preservation Approaches: These strategies maintain original parameters while updating new knowledge (Zheng et al., 2023; Jiang et al., 2024; Wang et al., 2024d; Hartvigsen et al., 2023; Li et al., 2025). The emergence of Large Multimodal Models (LMMs) has also advanced the field of multimodal knowledge editing. These works mainly focus on the editing of multimodal common knowledge (Cheng et al., 2023; Ma et al., 2024; Pan et al., 2024; Wang et al., 2024c) or multimodal factual knowledge (Li et al., 2024a; Zhang et al., 2024d; Huang et al., 2024b). However, the application of multimodal knowledge editing in the context of domain-specific knowledge, like autonomous driving remains underexplored.
5.2. Large Multi-modal Models for Autonomous Driving